OpenStudy (anonymous):

Jimmy's Chocolate shop sells two different fudge bars. The dark chocolate fudge bar sells for $6.00 per pound and the milk chocolate fudge bar sells for $4.00 per pound. You purchased a combination of these bars for all of your family and friends for the holidays and your bill came to $120. This scenario can be modeled by the equation 6d + 4m = 120. Which of the following is not a possible combination of dark chocolate bars and milk chocolate bars? Answer 12 dark bars and 12 milk bars 14 dark bars and 18 milk bars 10 dark bars and 15 milk bars 6 dark bars and 21 milk bars

OpenStudy (anonymous):

just plug in and see which ones dont work lol

OpenStudy (ash2326):

d= no. of dark chocolates m= no. of milk chocolates We have \[6d+4m=120\] @JWALKER295 we have 4 options. I'll check the last one and rest all you try 6 dark and 21 milk so d=6, m=21 \[6\times 6+4\times 12=120\] \[36+84=120\] \[120=120\] So this is a possible combination. Now you check the other options

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

i got is the second one thank you

OpenStudy (anonymous):

ur welcome!!!

OpenStudy (anonymous):

14 dark and 18 milk